OpenStudy (anonymous):

What is the length, c, of the hypotenuse of a right triangle? the width is 24 and the length is 10

OpenStudy (anonymous):

A^2 + b^2 = c^2 10^2 + 24^2 =c^2 100 + 576 = c^2 676 = c^2 square root to get final answer C = 26.

OpenStudy (anonymous):

thanks